# Graph-Based Approximate Nearest Neighbor Search Revisited: Theoretical Analysis and Optimization

### v2

@@
Abstract
Graph-based approaches to approximate nearest neighbor search (ANNS) have achieved remarkable success in enabling fast, high-recall retrieval on billion-scale vector datasets. Among them, the Sparse Neighborhood Graph (SNG) has emerged as a widely adopted graph structure due to its superior search performance. However, the theoretical understanding of SNG remains limited, leading to reliance on heuristic-based and often suboptimal truncation strategies. In this work, we aim to bridge the gap between theory and practice by providing formal guarantees for graph-based ANNS methods and proposing principled optimization strategies for the truncation parameter. By characterizing the index construction process through martingale-based analysis, we show that the degree of the index graph is $O(n^{2/3+\epsilon})$, where $\epsilon$ is an arbitrarily small constant. Furthermore, we prove that the expected search path length during query processing is $O(\log n)$. Based on these theoretical insights, we introduce a novel and principled method for selecting the truncation parameter $R$ in SNG. Experimental results demonstrate that our method achieves comparable or superior performance in terms of query latency and Recall@10 compared to commonly used binary search heuristics, while yielding 2x to 9x speedups in overall index construction.
1 Introduction
@@
NNS-ANNS Problems. Nearest neighbor search (NNS) is a fundamental problem in computer science KNN1 ; KNN2 and has become a fundamental research topic across various fields such as information retrieval IRTV ; image , machine learning ICDEML ; RAG and data mining Dataminingsig2017 . Formally, given a dataset $P=\{p_{1},\ldots,p_{n}\}\subset\mathbb{R}^{d}$ of cardinality $n$ and query point $q\in\mathbb{R}^{d}$, the NNS problem requires finding the $k$ closest points to $q$ under an appropriate distance metric $\|\cdot\|$. The computational complexity of exact NNS grows prohibitively with both dimensionality $d$ due to the well-documented curse of dimensionality Curse and the expanding dataset size $n$. This fundamental limitation has motivated the development of Approximate Nearest Neighbor Search (ANNS) algorithms, which achieve significant efficiency improvements while sacrificing minimal accuracy.
@@
Graph-Based ANNS. Modern ANNS approaches achieve remarkable efficiency-accuracy tradeoffs and can be categorized into four categories: tree-based approaches tree1 ; KD-TREE , hashing-based approaches LSHarshing ; hash2 ; hash3 ; hash4 , vector quantization-based approaches quantti1 ; RabitQ , and graph-based approaches Graph4 ; Graph3 ; Graph1 ; UNI1 . Among these, graph-based approaches have demonstrated particular success in practical applications Graph2024VLDB ; Graph1 , owing to their ability to capture local neighborhood structures. The graph-based ANNS framework operates through two sequential phases: an index construction phase that builds a graph preserving neighborhood relationships, followed by an query processing phase that employs the GreedySearch algorithm NSG19 ; NEURIPS2019 ; HNSW2020 to iteratively traverse the graph along paths until convergence to the query point. Numerous graph construction strategies have been proposed, each differing in how neighbors are selected and edges are formed.
Theoretical Limitations. Despite their empirical success, graph-based ANNS methods remain theoretically underexplored, primarily due to the stochastic nature of high-dimensional data and the combinatorial complexity of graph construction. Existing theoretical studies often rely on simplified assumptions and focus on analyzing key structural properties such as degree, query complexity, and search path length. For instance, sublinear query complexity has been established for threshold-based nearest neighbor graphs [^1] A threshold-based nearest neighbor graph connects two nodes if the distance between them is below a specified threshold value. under spherical data models ICML2020 , while navigable graphs [^2] A graph is navigable if the search algorithm can find a path from any starting vertex to any target vertex in the dataset. constructed under a binomial distribution exhibit an average degree of $O(n^{\beta})$ for $\beta<1/2$ Haya24 , although such assumptions are often unrealistic in practical settings. The Monotonic Search Network (MSNET) [^3] The Monotonic Search Network is a navigable graph where the path between any two nodes follows a monotonic trajectory. is known to achieve a logarithmic search path length NSG19 . The Sparse Neighborhood Graph (SNG) Arya93 is a variant designed to further reduce graph density while preserving search quality in practice. It has been broadly adopted in modern ANNS systems, including DiskANN NEURIPS2019 , HNSW HNSW2020 , and NSG NSG19 . The worst-case search analysis of SNG indyk2023worst relies on dataset aspect ratios [^4] The aspect ratio is the ratio of the maximum to minimum distance between any pair of points in the dataset., yielding bounds that are less interpretable compared to those based on cardinality.
@@
Practical Limitations and Parameter Tuning. To reduce index construction time, a common practice in SNG construction is to impose a maximum degree of node through a truncation parameter $R$. However, existing methods typically rely on parameter sweeping NEURIPS2019 , most commonly implemented as a binary search guided by the downstream search performance of the constructed graph. This tuning process for $R$ is computationally expensive, highlighting the need for a theoretically grounded approach to its selection—one that ensures high search performance while reducing tuning overhead.
1.1 Main Results
To address the dual challenges of theoretical understanding and practical parameter tuning in SNG-based approximate nearest neighbor search, this study develops a rigorous theoretical framework for graph construction under realistic random data assumptions, with a focus on the SNG. Building on this foundation, we further derive a marginal optimal principle-based strategy for parameter optimization. Specifically, we provide tighter bounds for both the indexing and query processes and offer an analytical expression for the key graph truncation parameter in SNG, enabling direct performance gains during index construction.
Degree Bound of the SNG Graph. We establish a tighter degree bound for SNG, improving upon the prior rough bound of $O(n)$ from NEURIPS2019 :
Theorem 1.
This result highlights the sparsity of the graph and the effectiveness of pruning during index construction. It implies that the SNG remains sparser than a complete graph. The proof idea and sketch are provided in Section 3, and the full proof is deferred to Appendix C.4.
Search Path Length on SNG. Building on the degree bound, we further demonstrate that the expected search path length in SNG is logarithmic in the dataset size:
Theorem 2.
The sketch of the proof is provided in Section 4, and the full argument appears in Appendix C.5.
Optimizing Parameter R. Based on these findings, we analyze the graph construction complexity under the truncated setting and derive an analytical expression for the optimal truncation parameter $R$, as presented in Section 5.2. This theoretically grounded formulation enables principled and efficient parameter selection, removing the need for manual tuning. Empirically, in the search phase, our optimized $R$ achieves comparable or improved latency and Recall@@10 compared to traditional binary search-based heuristics. In the index construction phase, our optimized approach delivers 2x to 9x acceleration across three standard benchmark datasets.
2 Preliminaries
2.1 Overview of SNG Construction Algorithm
The construction process of SNG proceeds as follows. Given a dataset $P=\{p_{1},\ldots,p_{n}\}\subset\mathbb{R}^{d}$, for each point $p\in P$, a candidate set $S=P\setminus\{p\}$ is initialized. While $S$ is non-empty, the algorithm selects the nearest neighbor $p^{*}=\arg\min_{s\in S}\|p-s\|$, adds a directed edge from $p$ to $p^{*}$, and prunes the candidate set by removing all points $p^{\prime}\in S$ that satisfy the condition $\|p-p^{\prime}\|>\|p^{*}-p^{\prime}\|$. In particular, under the truncated setting, the iterative process terminates after at most $R$ iterations.
To prevent pathological cases, such as points aligned on a straight line that could result in a graph diameter of $O(n)$, Vamana algorithm NEURIPS2019 introduces a relaxation parameter $\alpha$, modifying the pruning condition to $\|p-p^{\prime}\|>\alpha\cdot\|p^{*}-p^{\prime}\|$ [^5] In SNG constructions used in HNSWHNSW2020 and NSGNSG19 algorithm, the $\alpha$ parameter is not included and can be interpreted as being implicitly set to 1. The results of this paper also hold when incorporating parameter $\alpha$, for a comprehensive discussion, please refer to the end of Appendix C.4.. In practice, $\alpha$ is often set to a fixed value slightly greater than 1 and does not require tuning OOD .
2.2 Probabilistic Framework
Stochastic Modeling of the Pruning Process. To analyze the SNG construction, we establish the following probabilistic framework. Assume that $p$ is the point being processed. Let $t\in\{0,1,2,\dots\}$ index the iterations, and at iteration $t$, the candidate set is denoted $S_{t}$, initialized as $S_{0}=P\setminus\{p\}$. The corresponding processed points set is defined as $S^{\prime}_{t}=P\setminus(S_{t}\cup\{p\})$, i.e., $S^{\prime}_{t}$ contains all pruned points and previously selected nearest neighbors. In the non-truncated case, the process terminates when $S_{t}=\emptyset$, or equivalently, when $S^{\prime}_{t}=P\setminus\{p\}$. Due to the randomness of point distributions in high-dimensional space, the number of pruned points at each iteration is also stochastic, resulting in randomness in the graph structure. Consequently, the pruning process is inherently random. This correspondence between iteration counts and the evolution of $S_{t}$ and $S^{\prime}_{t}$ allows us to treat them as two stochastic processes. We formalize this with the following probabilistic setup. For definitions of foundational probability theory concepts, we refer readers to Appendix D.

### v3

# Sparse Neighborhood Graph-Based Approximate Nearest Neighbor Search Revisited: Theoretical Analysis and Optimization

@@
Abstract.
Graph-based approaches to approximate nearest neighbor search (ANNS) enable fast, high-recall retrieval on billion-scale vector datasets. Among them, the Sparse Neighborhood Graph (SNG) is widely used due to its strong search performance. However, the lack of theoretical understanding of SNG leads to expensive tuning of the truncation parameter that controls graph sparsification. In this work, we present OPT-SNG, a principled framework for analyzing and optimizing SNG construction. We introduce a martingale-based model of the pruning process that characterizes the stochastic evolution of candidate sets during graph construction. Using this framework, we prove that SNG has a maximum out-degree of $O(n^{2/3+\epsilon})$, where $\epsilon>0$ is an arbitrarily small constant, and an expected search path length of $O(\log n)$. Building on these insights, we derive a closed-form rule for selecting the optimal truncation parameter $R$, thereby eliminating the need for costly parameter sweeping. Extensive experiments on real-world datasets demonstrate that OPT-SNG achieves an average $5.9\times$ speedup in index construction time, with peak improvements reaching $15.4\times$, while consistently maintaining or improving search performance.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/maxila320/OPT-SNG.
1. Introduction
@@
Nearest Neighbor Search (NNS) is a fundamental problem in computer science (KNN1; KNN2; knn3; KNN4; knn5; knn6) and has become a fundamental research topic across various fields such as information retrieval (IRTV; image; retr2), data mining (Dataminingsig2017; datamining2; datamining3), and large language models (RAG; RAG2; ICDEML). Formally, given a dataset $P=\{p_{1},\ldots,p_{n}\}\subset\mathbb{R}^{d}$ of cardinality $n$ and query point $q\in\mathbb{R}^{d}$, the NNS problem requires finding the $k$ closest points to $q$ under an appropriate distance metric $\|\cdot\|$. The computational complexity of exact NNS grows prohibitively with both dimensionality $d$ due to the well-documented curse of dimensionality (Curse; curse2; curse3; curse4; curse5) and the expanding dataset size $n$. This fundamental limitation has motivated the development of Approximate Nearest Neighbor Search (ANNS) algorithms, which achieve significant efficiency improvements while sacrificing minimal accuracy.
@@
Modern ANNS approaches achieve remarkable search performance and fall into four categories: tree-based approaches (tree1; KD-TREE), hashing-based approaches (LSHarshing; hash2; hash3; hash4), vector quantization-based approaches (quantti1; RabitQ), and graph-based approaches (Graph4; Graph3; Graph1; UNI1). Among these methods, graph-based approaches have demonstrated particular success in practical applications (Graph2024VLDB; Graph1), as they effectively capture and exploit local neighborhood structures to enable navigation in high-dimensional spaces. A crucial observation is that despite the diversity of graph construction strategies—which vary in neighbor selection and edge formation rules—most state-of-the-art graph-based ANNS systems share a unified query processing paradigm. The graph-based ANNS framework operates through two complementary phases: (1) an index construction phase that builds a proximity graph preserving neighborhood relationships through various graph refinement strategies, and (2) a query processing phase that employs the GreedySearch algorithm (NSG19; NEURIPS2019; HNSW2020) to iteratively traverse the graph along decreasing-distance paths until convergence.
Sparse Neighborhood Graph (SNG) is a widely used variant designed to further reduce graph density while preserving practical search quality (Arya93). It has been widely adopted in several state-of-the-art systems, including DiskANN (NEURIPS2019), HNSW (HNSW2020) and NSG (NSG19). The worst-case search analysis of SNG by Indyk et al. (indyk2023worst) provides a framework based on aspect ratios [^1] The aspect ratio is the ratio of the maximum to minimum distance between any pair of points in the dataset.. Complementary to this, a developing and expected cardinality-based framework is both meaningful and necessary for understanding SNG behavior under random or empirical data distributions.
@@
The construction process of SNG typically involves two main phases: (1) Initial graph generation through a random or approximate k-NN graph, and (2) Graph refinement through iterative pruning(revisiting2025). The pruning phase is the computational bottleneck, with $O(n)$ worst-case complexity per node, as it requires scanning all candidate neighbors and evaluating geometric conditions that ensure the graph supports fast convergence of the GreedySearch algorithm. To reduce index construction time, a common practice in SNG construction is to impose a maximum node degree via a truncation parameter $R$. However, existing methods typically determine $R$ through parameter sweeping (or a parameter study) (NEURIPS2019), commonly implemented as a binary search guided by the downstream search performance of the constructed graph. This tuning process is computationally expensive, as each iteration requires constructing and evaluating a full index. These limitations highlight the need for a theoretically grounded approach to selecting RR—one that ensures high search performance while reducing tuning overhead.
To address both the theoretical and practical challenges in SNG-based approximate nearest neighbor search, we propose OPT-SNG, a martingale-based probabilistic framework for analyzing and optimizing SNG construction. In this framework, we characterize the SNG construction process as a discrete stochastic process driven by successive pruning operations. By exploiting the martingale structure naturally arising from the evolution of candidate sets, we establish that the expected trajectory can be characterized by an associated ordinary differential equation (ODE). Furthermore, we prove that the out-degree of each node corresponds precisely to the stopping time of this stochastic process. Within this framework, we derive bounds on two fundamental structural properties of SNGs—the maximum out-degree and the expected search path length. Building upon these theoretical results, we further quantify index construction complexity, which serves as a principled foundation for optimizing the truncation parameter $R$ in SNG-based ANNS systems.
We conclude our contributions as follows.
  • We introduce OPT-SNG, the first martingale-based probabilistic model of SNG pruning dynamics. By formulating the pruning process as a submartingale with respect to a natural filtration, we transform graph refinement into an analyzable probabilistic system. This enables the application of concentration inequalities and differential equation methods to study graph evolution.
  • We establish a degree bound of $O(n^{2/3+\epsilon})$ for any small constant $\epsilon>0$, substantially improving over the previous coarse $O(n)$ bound from (NEURIPS2019). This result highlights the inherent sparsity of SNGs and demonstrates that graph structure can be controlled through proper pruning dynamics analysis.
  • We prove that GreedySearch on SNGs converges in $O(\log n)$ expected steps. This analytical guarantee validates the intuition behind SNG’s empirical efficiency and provides theoretical justification.
  • Using the above theoretical foundations, we analyze the asymptotic complexity of the truncated SNG construction process and derive a closed-form expression for the optimal truncation parameter $R$. This yields a principled and automated parameter selection method, avoiding costly parameter sweeping over $R$. Empirically, our method achieves comparable or superior search performance while reducing index construction time by an average of $5.9\times$ across datasets, with peak speedups reaching $15.4\times$.
All detailed proofs of lemmas and theorems are provided in the full version (Ma2025) of this work on our GitHub repository.
The remainder of this paper is organized as follows. Section 2 reviews the necessary background on SNG-based ANNS, formally defines the problem setting, and introduces the key algorithms considered in this study. In Section 3, we present our probabilistic modeling framework that captures the stochastic dynamics of the pruning process. Sections 4 and 5 derive the main theoretical bounds on graph degree and search path length, respectively. Section 6 describes our optimized construction algorithm and the principled method for truncation parameter selection. Experimental results validating our theoretical predictions and demonstrating practical speedups are reported in Section 7. We survey the related works and position our contributions within the broader literature in Section 8, and Section 9 concludes with a summary and directions for future research.
2. $k$-ANN problem and Sparse Neighborhood Graphs
In this section, we formalize the approximate nearest search problem formulation, review existing SNG-based ANNS construction, and summarize the key algorithmic components relevant to our theoretical analysis. Given a dataset $P=\{p_{1},\ldots,p_{n}\}\subset\mathbb{R}^{d}$ of $n$ points in $d$-dimensional Euclidean space with distance metric $\|\cdot\|$, we consider a query point $q\in\mathbb{R}^{d}$. See Table 1 for a summary of key notations used throughout.
2.1. Problem Formulation
We formally define the $k$-approximate nearest neighbor search problem as follows.
Definition 0 ($k$-Approximate Nearest Neighbor Search).
Given a dataset $P\subset\mathbb{R}^{d}$, a query point $q\in\mathbb{R}^{d}$, and an integer $k\geq 1$, the $k$-approximate nearest neighbor search (K-ANNS) problem aims to return a set $A\subseteq P$ of $k$ data points that are closest to $q$, subject to a desired trade-off between recall and query latency.
Graph-based ANNS methods address this problem by modeling the dataset $P$ as the vertex set of a graph and constructing edges between data points to facilitate efficient search. Typically, the construction process begins with an initial graph, such as a random graph or a $k$-nearest neighbor graph, and then refines this structure through iterative edge pruning. Formally, these methods construct a graph $G=(P,E)$, where each vertex corresponds to a data point, and the edge set $E$ is designed to preserve neighborhood proximity, enabling fast approximate nearest neighbor search via local graph traversal.
2.2. GreedySearch Algorithm
Despite differences in index construction, many graph-based ANNS methods share a common query-processing phase: a greedy best-first traversal known as GreedySearch. Representative examples include KGraph (Kgraph), DiskANN (NEURIPS2019), NSG (NSG19), HNSW (HNSW2020), DPG (DPG), and SPTAG (SPANN2021). The procedure is described in Algorithm 1.
Starting from an entry point $s$, the algorithm maintains (1) a visited set and (2) a priority queue of candidate nodes ordered by distance to the query $q$. At each iteration, the closest unexpanded vertex $\hat{p}$ is extracted from the priority queue. Its outgoing neighbors are then explored and added to the priority queue if they have not been visited, and simultaneously, $\hat{p}$ is added to the visited set. The process continues until either the search width reaches the predefined budget $L$ or the priority queue becomes empty. The parameter $L$ controls the search width and governs the recall and latency trade-off: larger values of $L$ explore more candidate nodes, improving recall at the cost of increased search time.
Input : graph $G=(P,E)$, query point $q$, entry point $s$, $k$ for top-$k$, search width $L$
Output : top-$k$ approximate nearest neighbors of $q$
[…]
/* closest unexpanded point */
[…]
return the $k$ points in VisitedVisited closest to $q$
Algorithm 1 GreedySearch($G,q,s,k,L$)
13
13
13
13
13
13
13
13
13
13
13
13
13
Maintaining and accelerating the convergence of GreedySearch is a central objective in graph construction. In particular, ensuring graph connectivity while preserving monotonic search paths is both subtle and essential for guaranteeing fast convergence and high recall. The Sparse Neighborhood Graph (SNG) (Arya93) has been shown to effectively achieve these objectives by iteratively selecting nearest neighbors and pruning redundant candidates (indyk2023worst; Haya24).
2.3. Sparse Neighborhood Graph Construction
A core operation in SNG construction is the pruning procedure, which selectively retains informative neighbors while aggressively eliminating redundant connections. This procedure is a fundamental building block in several state-of-the-art graph-based ANNS systems, including DiskANN (NEURIPS2019), NSG (NSG19), and HNSW (HNSW2020).
[…]
Input: point $p$, candidate set $S$, pruning parameter $\alpha\geq 1$, degree bound $R$
Output: pruned neighbor set $N_{\text{out}}(p)$ with $|N_{\text{out}}(p)|\leq R$
[…]
21
21
21
21
21
21
21
21
21
21
21
21
21
21
21
21
21
21
21
21
21
The algorithm proceeds iteratively: in each iteration, it selects the nearest point $p^{*}$ from the candidate set $S$, adds $p^{*}$ to the neighbor set $N_{\text{out}}(p)$, and then prunes candidates from $S$. Specifically, a candidate $p^{\prime}$ is removed if it satisfies the blocking condition $\|p-p^{\prime}\|\geq\alpha\cdot\|p^{*}-p^{\prime}\|.$ This condition captures the intuition that $p^{\prime}$ is blocked by the recently selected neighbor $p^{*}$—that is, accessing $p^{\prime}$ through the intermediate neighbor $p^{*}$ is at least as efficient as a direct connection. Consequently, the direct edge $p\to p^{\prime}$ is redundant and can be safely removed without compromising graph navigability.
Two parameters control the pruning behavior:
Pruning Aggressiveness Parameter $\alpha$: In the non-relaxed setting ($\alpha=1$), the blocking condition is symmetric and purely geometric. However, to avoid pathological configurations (e.g., collinear points), the Vamana algorithm (NEURIPS2019) introduces $\alpha>1$, which relaxes the blocking condition and ensures bounded graph diameter. In practice, $\alpha$ is typically fixed in the range $1\leq\alpha\leq 2$ (OOD) and does not require extensive tuning, as the SNG structure remains robust across this range.
Degree Truncation Parameter $R$: To control graph density and construction cost, a maximum degree limit $R$ is imposed. The algorithm terminates after exactly $R$ neighbors have been selected, bounding the out-degree of each node. This truncation ensures that construction remains tractable while allowing a trade-off between graph quality and computational efficiency.
A crucial property of SNG-based construction is that it preserves monotonic searchability (NSG19). Specifically, for any point $p$, there exists at least one monotonic path consisting of edges that lead to progressively closer points, eventually reaching the nearest neighbor of $p$. This property is essential for the effectiveness of GreedySearch, as it guarantees that each expansion step can make consistent progress toward the target.
Despite its empirical success across multiple state-of-the-art systems (DiskANN, NSG, HNSW), the theoretical understanding of SNG construction remains incomplete. The introduction of relaxation parameter $\alpha$ substantially complicates the analysis: the blocking condition becomes asymmetric, and the geometric properties become more intricate. While Indyk et al. (indyk2023worst) provide worst-case analysis based on aspect ratios and Prokhorenkova and Shekhovtsov (Haya24) establish bounds under specific distributional assumptions, a comprehensive expected-case, cardinality-based framework characterizing SNG behavior has been lacking.
In particular, the truncation parameter $R$ remains a critical tuning knob. Existing practice relies on expensive parameter sweeping: practitioners repeatedly construct indices with different $R$ values, evaluate their search performance through binary or grid search, and select the value that optimizes the recall-latency trade-off. This iterative process is computationally prohibitive for large-scale datasets and obscures the fundamental principles governing the relationship between $R$, graph structure, and search efficiency.
Our OPT-SNG framework addresses these gaps by providing a principled, theoretically grounded approach to analyzing SNG dynamics and optimizing the truncation parameter $R$.
3. OPT-SNG Probabilistic Framework Setup
Figure 1. Illustration of the SNG process for a single point.
Stochastic Modeling of the Pruning Process. To analyze the construction of SNG from a dynamic perspective, we establish the following probabilistic framework that models the evolution of the pruning process. Consider a fixed point $p\in P$ being processed. Let $t\in\{0,1,2,\dots\}$ denote the iteration index, where each iteration is associated with two dynamically evolving sets.
Definition 0 (Candidate Set and Processed Set).
At iteration $t$, the candidate set is denoted by $S_{t}$. It is initialized as $S_{0}=P\setminus\{p\},$ meaning that, initially, the point $p$ is allowed to connect to all other points in the dataset. At each iteration, $S_{t}$ consists of the points that have not yet been pruned and have not been selected as neighbors. The corresponding processed set is defined as $S^{\prime}_{t}=P\setminus(S_{t}\cup\{p\}),$ which contains all points that have either been pruned or already selected as nearest neighbors in previous iterations.
In the non-truncated setting, the pruning process terminates when $S_{t}=\emptyset$, or equivalently when $S^{\prime}_{t}=P\setminus\{p\}$. Figure 1 provides a schematic illustration of this construction process. Initially ($t=0$), all other points belong to the candidate set $S_{0}$. At iteration $t=1$, the nearest neighbor $p_{1}^{*}$ is selected, and a subset of points $\Delta S_{1}$ is pruned and moved into the processed set $S^{\prime}_{1}$. This process repeats at subsequent iterations: at iteration $t$, a new nearest neighbor $p_{t}^{*}$ is selected from $S_{t-1}$, additional points are pruned, and the candidate set shrinks accordingly.
Due to the randomness of point distributions in high-dimensional space, the number of pruned points at each iteration is itself random, which induces randomness in the resulting graph structure. Consequently, the pruning procedure can be naturally viewed as a stochastic process. The correspondence between iteration indices and the evolution of $S_{t}$ and $S^{\prime}_{t}$ allows us to formalize the construction using tools from probability theory.